# Look-ahead Search on Top of Policy Networks in Imperfect Information Games

### 23 Dec 2023 v1

@@ 7. Conclusion
Policy-gradient algorithms proved to be powerful algorithms for solving huge problems in various settings. Recently, these algorithms were successfully applied to two-player zero-sum imperfect information games and outperformed human professionals in Stratego. However, the inability of these algorithms to incorporate search did not allow them to improve mistakes caused by network inaccuracies. We have presented an algorithm SePoT that can be used alongside any trajectory sampling policy-gradient algorithm in two-player zero-sum imperfect information games to train additional critic network without the need to perform search during training. We have shown that this network is sufficient to enable safe search during test time. The search we use differs from previous approaches since it is independent of the application of search in previous parts of the game, with the only limitation being the local size of the subgame. Thanks to this, SePoT can employ search even in larger games than previously possible while still having a choice to follow policy from the policy network.
The search in SePoT relies on the accurate reconstruction of the public state. Although a Constraint Satisfaction Problem Russell and Norvig (2009) can be utilized for games like Goofspiel, an efficient function to reconstruct public states remains an open problem for further investigation.
We introduced the notion of policy transformations, which allows training multiple expected values from just a single actor policy. The main limitation of SePoT stems from the fact that choosing the transformations affects search performance. A structured way to learn these transformations would further improve the performance of SePoT.
We showed how the multi-valued states connect to the value functions and how to use them as a usual value function, with just the added computation of a best response or by training the value function out of the multi-valued states. Furthermore, we have discussed how the multi-valued states could be expanded to allow both players to pick a strategy for the remainder of the game and how to use them with SePoT in this setting.
Our experimental results clearly demonstrate that the integration of search effectively reduces the exploitability compared to the trained policy network, even when using simple policy transformations. In head-to-head play, SePoT consistently outperforms RNaD with statistical significance in all tested game scenarios. We have shown that using the history critic to compute counterfactual values is possible. Still, in some games, it produces worse solutions than using these counterfactual values from previous resolves.
{acks}
@@
This research is supported by Czech Science Foundation (grant no. 18-27483Y). Computational resources were supplied by the project ”e-Infrastruktura CZ” (e-INFRA CZ LM2018140) supported by the Ministry of Education, Youth and Sports of the Czech Republic and also the OP VVV funded project CZ.02.1.01/0.0/0.0/16_019/0000765 “Research Center for Informatics” which are both gratefully ac- knowledged.
@@ Appendix A Proofs
This proof is heavily based on the proof of the original V-trace Espeholt et al. (2018), with the notable difference being that we discuss only the proof for finite horizon with discount factor $\gamma=1$. We can reformulate the V-trace operator as
$\mathcal{V}v_{i}(h)$ $=v_{i}(h)+\mathbb{E}_{\mu}\left[\sum^{l}_{t=0}\rho^{t}\Big{(}\prod_{0\leq q<t}c^{q}\Big{)}\Big{(}R_{i}^{t}+v_{i}(h^{t+1})-v_{i}(h^{t})\Big{)}\right]$
@@
$=$ $\sum^{l}_{t=0}\mathbb{E}_{\mu}\left[\Big{(}\prod_{0\leq q<t-1}c^{q}\Big{)}\rho^{t-1}\right]-\sum^{l}_{t=0}\mathbb{E}_{\mu}\left[\Big{(}\prod_{0\leq q<t}c^{q}\Big{)}\rho^{t}\right]+\rho^{-1}-\rho^{-1}$
$=$ $\sum^{l}_{t=0}\mathbb{E}_{\mu}\left[\Big{(}\prod_{0\leq q<t-1}c^{q}\Big{)}\rho^{t-1}\right]-\sum^{l+1}_{t=0}\mathbb{E}_{\mu}\left[\Big{(}\prod_{0\leq q<t-1}c^{q}\Big{)}\rho^{t-1}\right]+1$
@@
$=$ $\sum_{a_{i}\in\mathcal{A}_{i}(h^{t})}\pi_{\overline{\rho}}(s_{i}(h^{t}),a_{i})\Big{(}R_{i}^{t}+\sum_{h^{\prime}\in\mathcal{H}}p(h^{\prime}|h^{t},a_{i})v^{\pi_{\overline{\rho}}}_{i}(h^{\prime})-v^{\pi_{\overline{\rho}}}_{i}(h^{t})\Big{)}$
@@ B.1. Leduc hold’em
B.2. Imperfect Information Goofspiel
@@
In this work, we use two different versions of Goofspiel based on the shuffled order of the point cards. The first variation shuffles the point cards randomly. The second variation arranges the point cards from the highest value to the lowest, eliminating the chance player from the game. Original Goofspiel is played with thirteen cards. However, in this work, we also use Goofspiel, played with only five and eight private and point cards.
B.3. Battleships
@@
We have also used different versions of the Battleships, played on smaller boards, and each player has fewer ships.
Appendix C Additional experimental details
We used the same setting for all the hyperparameters for all the experiments. The setting is shown in Table 2. For the majority of hyperparameters, we adopted the settings established by the original authors in their work on RNaD Perolat et al. (2022). The differences are setting importance sampling clipping $\overline{\rho}$ to $\infty$ to ensure that the estimate of V-trace is unbiased. Furthermore, we changed the rate of updating the regularization policy, learning rate, batch size, and size of hidden layers to compensate for training on smaller games than the original authors did with Stratego. As a sampling policy $\mu$, we have used the actor policy, effectively avoiding importance sampling clipping by working on-policy.
@@ Appendix C Additional experimental details
batch size 128
hidden layers size 512, 512
CFR+ iterations in resolving 1500
Table 2. Parameters used during evaluation of SeRNaD
C.1. Policy Transformations
We have utilized a consistent set of 22 policy transformations across all the experiments on Goofspiel and Battleships. All of our transformations were independent of the infoset and kept throughout the whole game, meaning that, for example, the expected value of transformation 3 in a given history is if the opponent used this transformation in each decision point from this history onward. We introduce function $I:\mathcal{A}_{i}\rightarrow\mathbf{N}$ that for action $a_{i}$ returns the natural number. In Goofspiel, this corresponds to the card played. In Battleships, it corresponds to the row-major order of the tile on the grid. Additionally, we define $I^{R}(a_{i})$ as a complementary value, calculated as $I^{R}(a_{i})=\max_{a_{i}^{\prime}\in\mathcal{A}_{i}}I(a_{i}^{\prime})-I(a_{i})+1$. We split the transformations into multiple categories and elaborate on them.
  Identity transformation is required so that we can construct gadget game
  1. (1)
    $\pi^{\prime}(s_{i},a_{i})=\pi(s_{i},a_{i})$
Uniform policy
  1. (2)
    $\pi^{\prime}(s_{i},a_{i})=\frac{1}{|\mathcal{A}_{i}(s_{i})|}$
Transformation that increases probability of playing actions that are played with small probability in the original policy
  1. (3)
    $\pi(s_{i},a)=\pi(s_{i},a)$
Increasing the probability of playing actions with higher value of $I$
  1. (4)
    $\pi(s_{i},a)=I(a)\pi(s_{i},a)$
  2. (5)
    $\pi(s_{i},a)=I(a)^{2}\pi(s_{i},a)$
  3. (6)
    $\pi(s_{i},a)=e^{I(a)}\pi(s_{i},a)$
Increasing the probability of playing actions with higher value of $I^{R}$
  1. (7)
    $\pi(s_{i},a)=I^{R}(a)\pi(s_{i},a)$
  2. (8)
    $\pi(s_{i},a)=I^{R}(a)^{2}\pi(s_{i},a)$
  3. (9)
    $\pi(s_{i},a)=e^{I^{R}(a)}\pi(s_{i},a)$
Increasing the probability of playing actions that have both high values of $I$ and $I^{R}$
  1. (10)
    $\pi(s_{i},a)=\min\Big{(}I(a)\pi(s_{i},a),I^{R}(a)\pi(s_{i},a)\Big{)}$
  2. (11)
    $\pi(s_{i},a)=\min\Big{(}I(a)^{2}\pi(s_{i},a),I^{R}(a)^{2}\pi(s_{i},a)\Big{)}$
  3. (12)
    $\pi(s_{i},a)=\min\Big{(}e^{I(a)}\pi(s_{i},a),e^{I^{R}(a)}\pi(s_{i},a)\Big{)}$
Increasing the probability of playing actions that either have high value of $I$ or $I^{R}$
  1. (13)
    $\pi(s_{i},a)=\max\Big{(}I(a)\pi(s_{i},a),I^{R}(a)\pi(s_{i},a)\Big{)}$
  2. (14)
    $\pi(s_{i},a)=\max\Big{(}I(a)^{2}\pi(s_{i},a),I^{R}(a)^{2}\pi(s_{i},a)\Big{)}$
  3. (15)
    $\pi(s_{i},a)=\max\Big{(}e^{I(a)}\pi(s_{i},a),e^{I^{R}(a)}\pi(s_{i},a)\Big{)}$
Removing some actions from the policy
  1. (16)
[…]
  2. (17)
[…]
  3. (18)
[…]
  4. (19)
[…]
  5. (20)
[…]
  6. (21)
[…]
  7. (22)
[…]
Appendix D Additional experimental results
In this section we show some additional results from experiments in Section 5.2 and 5.3. All of these experiments are performed on small Goofspiel and confirms the results in Section 5.
D.1. Exploitability in smaller games
Figure 5. Exploitability in Goofspiel 5 with descending point card ordering from perspective of each player
(a) Player 1
(b) Player 2
Exploitability in Goofspiel 5 with descending point card ordering from perspective of each player
Figure 5. Exploitability in Goofspiel 5 with descending point card ordering from perspective of each player
Figure 6. Exploitability in Goofspiel 5 with randomized point card ordering from perspective of each player
(a) Player 1
(b) Player 2
Exploitability in Goofspiel 5 with randomized point card ordering from perspective of each player
Figure 6. Exploitability in Goofspiel 5 with randomized point card ordering from perspective of each player
We further examine the exploitability in small Goofspiel. In Figure 5, we show how the exploitability changes with additional training iterations in Goofspiel 5 with descending point card ordering with roughly 18 thousand unique histories. In Figure 6, we show the same experiment in Goofspiel 5 with randomized point card ordering. In both of these experiments, we see that using look-ahead 2 and 3 has roughly the same results, which are better than look-ahead 1 or playing according to the policy network.
D.2. Counterfactual values in gadget games
Figure 7. Exploitability in Goofspiel 5 with descending point card ordering from perspective of each player with look-ahead limit 1.
(a) Player 1
(b) Player 2
Exploitability in Goofspiel 5 with descending point card ordering from perspective of each player with look-ahead limit 1.
Figure 7. Exploitability in Goofspiel 5 with descending point card ordering from perspective of each player with look-ahead limit 1.
Figure 8. Exploitability in Goofspiel 5 with descending point card ordering from perspective of each player with look-ahead limit 2.
(a) Player 1
(b) Player 2
Exploitability in Goofspiel 5 with descending point card ordering from perspective of each player with look-ahead limit 2.
Figure 8. Exploitability in Goofspiel 5 with descending point card ordering from perspective of each player with look-ahead limit 2.
Figure 9. Exploitability in Goofspiel 5 with descending point card ordering from perspective of each player with look-ahead limit 3.
(a) Player 1
(b) Player 2
Exploitability in Goofspiel 5 with descending point card ordering from perspective of each player with look-ahead limit 3.
Figure 9. Exploitability in Goofspiel 5 with descending point card ordering from perspective of each player with look-ahead limit 3.
Figure 10. Exploitability in Goofspiel 5 with randomized point card ordering from perspective of each player with look-ahead limit 1.
(a) Player 1
(b) Player 2
Exploitability in Goofspiel 5 with randomized point card ordering from perspective of each player with look-ahead limit 1.
Figure 10. Exploitability in Goofspiel 5 with randomized point card ordering from perspective of each player with look-ahead limit 1.
Figure 11. Exploitability in Goofspiel 5 with randomized point card ordering from perspective of each player with look-ahead limit 2.
(a) Player 1
(b) Player 2
Exploitability in Goofspiel 5 with randomized point card ordering from perspective of each player with look-ahead limit 2.
Figure 11. Exploitability in Goofspiel 5 with randomized point card ordering from perspective of each player with look-ahead limit 2.
Figure 12. Exploitability in Goofspiel 5 with randomized point card ordering from perspective of each player with look-ahead limit 3.
(a) Player 1
(b) Player 2
Exploitability in Goofspiel 5 with randomized point card ordering from perspective of each player with look-ahead limit 3.
Figure 12. Exploitability in Goofspiel 5 with randomized point card ordering from perspective of each player with look-ahead limit 3.
In this section, we present additional results that compare using the critic network to compute counterfactual values or those from previous resolves within continual resolution. In Figure 7-9 we compare the effect of using these counterfactual values in Goofspiel 5 with descending point card ordering for different look-ahead limits. Results from the same experiment in Goofspiel 5 with randomized point card ordering are shown in Figure 10-12. In all of these Goofpsiel experiments, using the history critic to compute the counterfactual values yields the only slightly worse solutions than reusing them from previous searches.

### 06 Feb 2024 v2

@@ 7 Conclusion
Policy-gradient algorithms proved to be powerful algorithms for solving huge problems in various settings. Recently, these algorithms were successfully applied to two-player zero-sum imperfect information games and outperformed human professionals in Stratego. However, these algorithms could not incorporate search to improve mistakes caused by network inaccuracies. We have presented an algorithm SePoT that can be used alongside any trajectory sampling policy-gradient algorithm in two-player zero-sum imperfect information games to train additional critic network without any search during training. We have shown that this network is sufficient to enable safe search during test time. The search we use differs from previous approaches since it is independent of the application of search in previous parts of the game, with the only limitation being the local size of the subgame. Thanks to this, SePoT can employ search even in larger games while still having a choice to follow policy from the policy network. Our experimental results demonstrate that SePoT is able to outperform RNaD in most of the tested scenarios.
Acknowledgments
@@
This research is supported by Czech Science Foundation (grant no. 18-27483Y) and Grant Agency of the Czech Technical University in Prague, grant No. SGS22/168/OHK3/3T/13. Computational resources were supplied by the project ”e-Infrastruktura CZ” (e-INFRA CZ LM2018140) supported by the Ministry of Education, Youth and Sports of the Czech Republic and also the OP VVV funded project CZ.02.1.01/0.0/0.0/16_019/0000765 “Research Center for Informatics” which are both gratefully acknowledged.
@@ Appendix A Proofs
This proof is heavily based on the proof of the original V-trace [Espeholt et al., 2018], with the notable difference being that we discuss only the proof for finite horizon with discount factor $\gamma=1$. We can reformulate the V-trace operator as
$\mathcal{V}v_{i}$ $(h)=v_{i}(h)$
$+\mathbb{E}_{\mu}\left[\sum^{l}_{t=0}\rho^{t}\Big{(}\prod_{0\leq q<t}c^{q}\Big{)}\Big{(}R_{i}^{t}+v_{i}(h^{t+1})-v_{i}(h^{t})\Big{)}\right]$
@@
$=$ $\sum^{l}_{t=0}\mathbb{E}_{\mu}\left[\Big{(}\prod_{0\leq q<t-1}c^{q}\Big{)}\rho^{t-1}\right]$
$-\sum^{l}_{t=0}\mathbb{E}_{\mu}\left[\Big{(}\prod_{0\leq q<t}c^{q}\Big{)}\rho^{t}\right]+\rho^{-1}-\rho^{-1}$
$=$ $\sum^{l}_{t=0}\mathbb{E}_{\mu}\left[\Big{(}\prod_{0\leq q<t-1}c^{q}\Big{)}\rho^{t-1}\right]$
$-\sum^{l+1}_{t=0}\mathbb{E}_{\mu}\left[\Big{(}\prod_{0\leq q<t-1}c^{q}\Big{)}\rho^{t-1}\right]+1$
@@
$=$ $\sum_{a_{i}\in\mathcal{A}_{i}(h^{t})}\pi_{\overline{\rho}}(s_{i}(h^{t}),a_{i})$
@@ B.1 Leduc hold’em
B.2 Imperfect Information Goofspiel
@@
In this work, we use two different versions of Goofspiel based on the shuffled order of the point cards. The first variation shuffles the point cards randomly. The second variation arranges the point cards from the highest value to the lowest, eliminating the chance player from the game, making the game much smaller. Original Goofspiel is played with thirteen cards. However, in this work, we also use Goofspiel, played with only five and eight private and point cards.
B.3 Battleships
@@
We have used different versions of the Battleships, played on smaller boards, and each player has fewer ships.
Appendix C Additional experimental results
C.1 Counterfactual values in gadget games
This experiment aims to show that using the counterfactual values computed from history critic does not significantly worsen the solution compared to using them from previous resolves. We trained ten critic networks with the same hyperparameter setting for Leduc hold’em and Goofspiel with five cards and both descending point card ordering. Then, we solved each subgame separately, either by computing the counterfactual values for the gadget game from the history critic or reusing them from previous searches. In Leduc hold’em, we split the game before dealing a public card and after. In Goofspiel, each subgame had a depth limit of 1, so the value function is called after a single action by each player from the root of the subgame. The action in the gadget game is not counted towards the depth limit. We then combine these subgame policies into a policy for a whole game and compute the exploitability.
The results are shown in Figure 4, where we show both approaches’ mean and the 95% confidence interval. Using the counterfactual values from previous searches only slightly outperforms using the history critic. The results confirm that using critic network affects the quality of the solution in a minor way. With this insight, the remainder of our experiments were conducted with the computation of the counterfactual values from the critic network.
(a) Leduc hold’em
(b) Goofspiel 5 descending
Figure 4: Exploitability based on the training iterations, when using the counterfactual values from previous searches or computing them from history critic.
C.2 Predefined transformations
We have created 22 predefined transformations for Goofspiel. These transformations correspond to some strategies in Goofspiel that are strong in certain scenarios, like playing more higher value cards. We have then trained ten networks with the same trajectories. For each 500 iterations, we compared the exploitability when having a search depth-limit one between these predefined transformations and the neural network transformations. The results for Goofspiel 5 with descending and randomized order of point cards are shown in Figure 5.
The predefined transformations perform better in the tested scenario. This is to be expected, since the predefined transformations were specifically selected to perform well for Goofspiel. Therefore, SePoT performance could be further improved by specyfing domain-specific transformations, or by improving the automatic transformations.
(a) Goofspiel 5 descending
(b) Goofspiel 5 randomized
Figure 5: Exploitability in Goofspiel based on the training iterations, comparing predefined transformations with the neural network transformations
Appendix D Additional experimental details
We used the same setting for all the hyperparameters for all the experiments. The setting is shown in Table 2. For the majority of hyperparameters, we adopted the settings established by the original authors in their work on RNaD [Perolat et al., 2022]. The differences are setting importance sampling clipping $\overline{\rho}$ to $\infty$ to ensure that the estimate of V-trace is unbiased. Furthermore, we changed the rate of updating the regularization policy, learning rate, batch size, and size of hidden layers to compensate for training on smaller games than the original authors did with Stratego. As a sampling policy $\mu$, we have used the actor policy, effectively avoiding importance sampling clipping by working on-policy.
The only exception with different hyperparameter settings was Battleship 7x7, for which we have set the importance sampling clipping $\overline{\rho}=3$.
@@ Appendix D Additional experimental details
batch size 64
hidden layers size 1024, 1024
CFR+ iterations in resolving 800
Policy transformations 10
Table 2: Parameters used during evaluation of SeRNaD
In Table 3, we provide approximate amount of unique histories for each of the used games during our experiments
Game Histories
Goofspiel 5 descending points $1.8\cdot 10^{4}$
Goofspiel 5 randomized points $2.4\cdot 10^{6}$
Goofspiel 8 randomized points $9.1\cdot 10^{13}$
Goofspiel 13 descending points $4.9\cdot 10^{19}$
Goofspiel 13 randomized points $3.4\cdot 10^{28}$
Leduc hold’em $9.4\cdot 10^{3}$
Battleship 3x3 $\approx 10^{14}$
Battleship 5x5 $\approx 10^{57}$
Battleship 7x7 $\approx 10^{138}$
Table 3: Approximate amount of histories in tested games
Appendix E Pseudocodes
To improve the understandability of the full algorithm, we provide pseudocodes for training SePoT with some arbitrary trajectory-sampling-based policy-gradient algorithm and for gameplaying with CFR+. It is important to note that within CFR+, which performs a search, the counterfactual values have to be computed via the value function after the depth limit. This consists of retrieving multi-valued states for each history in the leaves and then computing the opponents’ best response value for each infosets of each player.
Algorithm 1 SePoT training
1:Sampling Policy $\mu$, Iterations $T$
2:$\theta^{\pi},\theta^{t},\theta^{v}\leftarrow$ InitializeNetworks()
3:for $t\leftarrow 1\text{ to }T$ do
4:     $\tau\leftarrow$ SampleTrajectories($\mu$)
5:     $\pi_{\text{old}}\leftarrow$ RetrievePolicyFromNetwork($\tau,\theta^{\pi}$)
6:     $\theta^{\pi}\leftarrow$ PolicyGradientUpdate($\tau,\mu,\theta^{\pi}$)
7:     $\pi_{\text{new}}\leftarrow$ RetrievePolicyFromNetwork($\tau,\theta^{\pi}$)
8:     $d\leftarrow\frac{\pi_{\text{new}}-\pi_{\text{old}}}{||\pi_{\text{new}}-\pi_{\text{old}}||}$
9:     $\theta^{t}\leftarrow$ UpdateTransformationNetwork($\tau,d,\theta^{t}$)
10:     $\pi^{t}\leftarrow$ ApplyTransformations($\pi,\theta^{t}$ )
11:     $\theta^{v}\leftarrow$ UpdateCriticNetwork($\tau,\mu,\pi^{t},\theta^{v}$)
12:end for
13:return $\theta^{\pi},\theta^{v}$
Algorithm 2 SePoT gameplay
1:Actor Weights $\theta^{\pi}$, Critic Weights $\theta^{v}$, Player $i$ Current decision node $s_{i}$
2:if CanPublicStateBeReconstructed($s_{i}$then
3:     $s_{0}\leftarrow$ ReconstructPublicState($s_{i}$)
4:     $r_{i},r_{c}\leftarrow$ RetrieveRanges($s_{0},\theta^{\pi}$)
5:     $u\leftarrow$ RetrieveMultiValuedStates($s_{0},\theta^{v}$)
6:     $v_{o}\leftarrow$ ComputeOpponentsRootCFVs($u,r_{i},r_{c}$)
7:     $\mathcal{G}^{G}\leftarrow$ ConstructGadgetGame($r_{i},r_{c},v_{o}$)
8:     $\pi\leftarrow$ CounterfactualRegretMinimizationPlus($\mathcal{G}^{G},\theta^{v}$)
9:else
10:     $\pi\leftarrow$ RetrievePolicyFromNetwork($s,\theta^{\pi}$)
11:end if
12:return $\pi$